Two students are 350 m apart and begin walking toward one another at constant rates. One travels at 1.6 m/s and other at 1.9 m/s. how long will they walk until they meet? how far has each of them gone?
Answer by mananth(16949)

let one travel x meters before they meet.
the other will walk 350-x meters
speed of one is 1.6 m / s
other is 1.9 m/s
..
the time they walk before they meet is the same.
x/1.6 = 350-x/1.9
1.9x=1.6(350-x)
1.9x=560-1.6x
2.5x=560
x=560/2.5
x= 224 m
the other travels 350-244=106 m
..
the time they meet will be
224/1.6 =140 seconds after they start